OpenStudy (anonymous):

Need help,....can someone explain about polarization and resolution?...got a test on it tomorow...formula for that too please..

OpenStudy (anonymous):

medal for the best explanation

OpenStudy (anonymous):

Polarized electromagnetic waves have electric vector oscillating in one direction direction in the plane perpendicular to the the direction of propagation, and they have the magnetic vector perpendicular to the electric. Usually, light is a mixture of many polarizations, but passing it through a polarizer "filters out" the other directions, giving polarized light. Light refracts differently depending on its polarization and scatters differently, also.

OpenStudy (anonymous):

Malus' Law: \[I =I _{\max}\cos ^{2}\theta\] Brewster's Law: \[\tan \theta=\frac{ n_{b} }{ n_{a} }\] I got those formulas from Sears Zemansky University Physics with Modern Physics

OpenStudy (anonymous):

charliealbert can you explain further about brewster's law and malus'

OpenStudy (anonymous):

Malus: intensity penetrating through polarizers crossed at angle @ or through one polarizer if the incident light is already polarized at angle @ to it. Brewster: incidence angle for complete polarization of reflected light.

OpenStudy (anonymous):

Resolution is a property of microscopes and telescope which allow the viewer to see two closely space objects as separate objects. Diffraction is the physical phenomenon that causes blurring around the edges of objects. If two objects are too close together this blurring can cause the two objects to appear as one object . Thus the objects are not resolved. The higher the resolution the sharper the images appear and the more fine detail can be seen. The human eye can resolve, i.e. see closely space objects, about 0.1 mm apart.at normal viewing distance.

OpenStudy (anonymous):

owh i see...thanks